My son is a student at the Virginia School of Deaf and has brought home 5 pages of homework printed from a text book.I am not able to provide you with the book name.
I am very frustrated with the problems, primarily due to the fact I haven't done this stuff for 30 years. How do I even attempt him with this stuff? Even the most basic I dont recall. Please help with one of the problems . maybe I can get a grasp of this with your assistamce.
-8+2x+5+11x

Found 2 solutions by checkley71, josmiceli:
Answer by checkley71(8403)

8+2X+5+11X=0
13X+13=0
13X=-13
X=-13/13
X=-1
PROOF
8+2*-1+5+11*-1=0
8-2+5-11=0
13-13=0
0=0

Answer by josmiceli(6793)

In order to be a problem with a solution you've got to have an equation
something = something else
I can simplify what you have, though
-8 + 2x + 5 + 11x
group the numbers and group the unknowns
-8 + 5 + 2x + 11x
perform additions of like terms
-3 + 13x
That's the answer if you just need to simplify